OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -6 and 162, respectively. Please explain this to me very carefully.

OpenStudy (campbell_st):

a term in a geometric series is \[T_{n} = ar^{n -1}\] so you have for the 2nd term \[-6 =ar\] and the 5th term \[162 = ar^4\] this can be written as \[162 = ar \times r^3\] substitute the information from the 2nd term \[162 = -6 \times r^3\] you can now solve for the common ration r when you have it, you will be able to solve for the 1st term a. hope this helps

OpenStudy (anonymous):

Why did we put 162 = ar * r^(3) ?

OpenStudy (anonymous):

because we know that \(-6 = ar\), so if we substitute it in we will be left with one variable \(r\) (the common ratio), and we can solve for \(r\) Then after you find \(r\) you can substitute it into \(-6 = ar\) to find \(a\)

OpenStudy (campbell_st):

because you are then able to find the common ratio... you get \[-27 = r^3\] solve for r and you'll then find the 1st term is a = 2

OpenStudy (anonymous):

what is the explicit rule?